OpenStudy (tylerballardwildcat):

What is the fourth term of (x + y)8? medal

OpenStudy (niksbhalla14mar):

did you get an answer to your earlier question. Please show me too..

OpenStudy (welshfella):

use the same formula in the last question.

OpenStudy (niksbhalla14mar):

Please tell me also.. i couldnt see that question's solution..

OpenStudy (welshfella):

the term will include the value of 8C3 The number of combinations of 3 from 8. Can you work that out?

OpenStudy (tylerballardwildcat):

im not good with this

OpenStudy (tylerballardwildcat):

im trying to catch on

OpenStudy (niksbhalla14mar):

How is What is the fourth term of (x + y)8? a question of combination. dint understand it al all... please explain

OpenStudy (owen_turner):

8/3 x^8-3 y^3 = 8!/3! (8-3) x^5 x^3 = 8x7x 6/3x2 x^5 x^3 = 56x^5 y^3

OpenStudy (welshfella):

Ah! I cant remember the details - I'll try and find you a link.

OpenStudy (welshfella):

Owen Turner has the answer

OpenStudy (niksbhalla14mar):

What did Owen Turner you did ? Have no idea at all...

OpenStudy (owen_turner):

Yes.

OpenStudy (welshfella):

https://en.wikipedia.org/wiki/Binomial_theorem

OpenStudy (welshfella):

the formula for the (r + 1)th term of (a + b)^n is nCr a^(n-r) b^r